OpenStudy (anonymous):

x^2 y^-3 z^3 ______________?? simpliy x^4 y^2 z^4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Remember that when you have to values with the same base, and you divide them, it's equivalent to subtracting the exponents. First, we can break it up like this: \[\frac{ x^2 y^{-3} z^3 }{ x^4 y^2 z^4 }=\frac{ x^2 }{ x^4 } + \frac{ y^{-3} }{ y^2 } + \frac{ z^3 }{ z^4 }\]

OpenStudy (anonymous):

Sorry those plus signs should be multiplication signs between the fractions

OpenStudy (anonymous):

But now, you can go piece by piece, remembering that dividing can be done by subtracting the exponents

OpenStudy (anonymous):

look at the file

OpenStudy (anonymous):

so whats the answer @vinnv226

OpenStudy (anonymous):

Go piece by piece. I'll show you how to do the x piece. Since we have the same base we can subtract the exponents: \[\frac{ x^2 }{ x^4 } = x^{2-4}=x^{-2}\] Now do the same thing with the other pieces

OpenStudy (anonymous):

y^-1 z^-1?? @vinnv226

OpenStudy (anonymous):

The z piece is right but the y piece isn't. It should be y^(-3 -2)=?

OpenStudy (anonymous):

y^-6

OpenStudy (anonymous):

@vinnv226??

OpenStudy (anonymous):

he is not coming back

OpenStudy (anonymous):

well is this correct @smartguy1124

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

-3 -2 = -5, not -6

OpenStudy (anonymous):

its y^-4

OpenStudy (anonymous):

but u were close @broderick365

OpenStudy (anonymous):

y^(-3) / y^(2) is going to be y^(-5), not y^(-4)